### v1

# Unsupervised Anomaly Detection through Mass Repulsing Optimal Transport

Abstract
Detecting anomalies in datasets is a longstanding problem in machine learning. In this context, anomalies are defined as a sample that significantly deviates from the remaining data. Meanwhile, optimal transport (OT) is a field of mathematics concerned with the transportation, between two probability measures, at least effort. In classical OT, the optimal transportation strategy of a measure to itself is the identity. In this paper, we tackle anomaly detection by forcing samples to displace its mass, while keeping the least effort objective. We call this new transportation problem Mass Repulsing Optimal Transport (MROT). Naturally, samples lying in low density regions of space will be forced to displace mass very far, incurring a higher transportation cost. We use these concepts to design a new anomaly score. Through a series of experiments in existing benchmarks, and fault detection problems, we show that our algorithm improves over existing methods.
Keywords Optimal Transport  $\cdot$ Anomaly Detection  $\cdot$ Fault Detection  $\cdot$ Cross Domain Fault Detection
1 Introduction
An anomaly, or an outlier, is a data point that is significantly different from the remaining data (Aggarwal, 2017), to such an extent that it was likely generated by a different mechanism (Hawkins, 1980). From the perspective of machine learning, Anomaly Detection (AD) wants to determine, from a set of examples, which ones are likely anomalies, typically through a score. This problem finds applications in many different fields, such as medicine Salem et al. (2013), cyber-security Siddiqui et al. (2019), and system monitoring Isermann (2006), to name a few. As reviewed in Han et al. (2022), existing techniques for AD are usually divided into unsupervised, semi-supervised and supervised approaches, with an increasing need for labeled data. In this paper, we focus on unsupervised AD, which does not need further labeling effort in constituting datasets. As discussed in Livernoche et al. (2024), the growing number of applications involving high-dimensional and complex data begs the need for non-parametric algorithms.
Meanwhile, Optimal Transport (OT) is a field of mathematics concerned with the transportation of masses at least effort Villani et al. (2009). In its modern treatment, one can conceive transportation problems between probability distributions, which has made an important impact in machine learning research (Montesuma et al., 2024a). Hence, OT is an appealing tool, as it can be estimated non-parametrically from samples from probability distributions. Likewise, the plethora of computational tools for computing OT (Peyré and Cuturi, 2020; Flamary et al., 2021) further stresses its usability.
In this context, the application of OT for AD is not straightforward, as we are interested in analyzing a single probability distribution, rather than 2. This paper puts forth a new OT problem between a distribution and itself, by restricting where a sample can send its mass to. More specifically, we design an exclusion zone, prohibiting samples from keeping its mass, or sending its mass to a small vicinity. Especially, we assume that anomalies lie in low density regions of space, with only a few samples in their vicinity. By restricting the transport of mass in the vicinity of samples, anomalies are naturally forced to send its mass to the high density region, which is assumed to be far away from the anomaly samples. Hence, anomalies will make a larger transportation effort than normal samples, which can find nearby samples outside the exclusion zone. We show a conceptual illustration of our method in Figure.
While OT has been previously used to compare and aggregate signals in the context of AD (Alaoui-Belghiti et al., 2019, 2020), to the best of our knowledge ours is the first general purpose OT-based algorithm for AD. Furthermore, we propose a new OT problem based on the engineering of the ground-cost, which has links to OT with repulsive costs (Di Marino et al., 2017). We benchmark our algorithm in a comprehensive list of datasets, including tabular, computer vision, and natural language processing proposed by Han et al. (2022), besides fault detection (Reinartz et al., 2021; Montesuma et al., 2024b).
This paper is organized as follows. Section 2 discusses related work in AD and OT. Section 3 discusses our proposed method, called Mass Repulsing Optimal Transport (MROT). Section 4 covers our experiments. Finally, section 5 concludes this paper.
2 Related Work
Anomaly Detection. Following Han et al. (2022), AD methods can be mainly divided into 3 categories. First, supervised methods consider AD through the lens of binary classification under class imbalance. Second, semi-supervised methods either consider partially labeled data, or labeled normal samples, so that an algorithm can be characterize what a normal sample is. The third, more challenging category is unsupervised AD, where the training data contains both anomalies and normal samples and labels are not available. This paper considers precisely the last setting. Next, we review ideas in unsupervised AD.
The first kind of methods rely on encoder-decode architectures to detect anomalies. The insight is that, by embedding data in a lower dimensional space, anomalies can be detected via the reconstruction error of the auto-encoding function. This is the principle of Principal Component Analysis (PCA) Shyu et al. (2003), which employs linear encoding and decoding functions, but also of kernelized versions Schölkopf et al. (1997); Hoffmann (2007), as well as neural nets Vincent et al. (2008); Bengio et al. (2013), which rely on non-linear embedding techniques.
The second type of strategies are based on the paradigm of 1-class classification. As Schölkopf et al. (1999) puts, the idea is to define a function that outputs 00 on a small, dense region of the space where normal samples lie, and $1$ elsewhere. In this context, Schölkopf et al. (1997) extends the celebrated Support Vector Machine (SVM) to AD, and Liu et al. (2008) extends Random Forests (RFs) of Breiman (2001).
A third kind of approaches focuses on neighborhoods and clustering, to model the data underlying probability distribution, especially through the density of samples over the space. This is the case of $k-$Nearest Neighbors ($k-$NN) (Ramaswamy et al., 2000), who use distances and nearest neighbors to determine anomalies, Local Outlier Factor (LOF) Breunig et al. (2000), who devised a score that measures the local deviation of a sample with respect its neighbors. Finally, Clustering-based LOF (CBLOF) He et al. (2003) proposed an extension of LOF based on the relative sizes of clusters within the data.
A fourth kind of emerging ideas in the field considers generative modeling for detecting anomalies. In this context, one devises a model for the data density $P$, for instance, through neural nets Goodfellow et al. (2014); Kingma and Welling (2013). The idea from these models, is that, either the data density can be accurately estimated (Dias et al., 2020), or some score can be derived from the generative model (Livernoche et al., 2023).
As we cover in the next section, our method uses nearest neighbors and OT to model, non-parametrically, the density of samples over the space. More specifically, we prohibit samples in OT to keep their mass, or sending it over a region of space defined through their $k-$NN. Differently from Ramaswamy et al. (2000) and Breunig et al. (2000), we do not rely on distances, which might not have a meaning in high-dimensions. Rather, we rely on the effort of transportation, measured through the samples’ mass times the ground-cost.
Optimal Transport with Repulsive Costs. In general OT theory (see, e.g., Section 3.1 below), samples are transported based on a ground-cost that measures how expensive it is to move masses between measures. In its original conception by Monge (1781) and Kantorovich (1942), this ground cost is the Euclidean distance $c(\mathbf{x}_{1},\mathbf{x}_{2})=\lVert\mathbf{x}_{1}-\mathbf{x}_{2}\rVert_{2}^{2}$. As reviewed in Di Marino et al. (2017), it may be interesting to consider repulsive costs, i.e. functions $c$ that are big when $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ are close to each other and small otherwise. An example of such costs, arising from physics, is the Coulomb interaction $c(\mathbf{x}_{1},\mathbf{x}_{2})=(\lVert\mathbf{x}_{1}-\mathbf{x}_{2}\rVert_{2})^{-1}$. Still following Di Marino et al. (2017), these kinds of transportation problems proved useful in physics, e.g., for quantum mechanics and $N-$body systems.
In this paper, we consider a different kind of repulsive cost, which we call the mass repulsive cost (see, e.g., Section 3.2 below). Our notion of cost defines an exclusion zone where sending mass is too costly. This exclusion zone is defined on the nearest neighbor analysis of points in a probability measure. As a result, our approach captures the local characteristics of the probability measure being analyzed, especially its density. A special characteristic of our approach is to give a sense of the transportation from a measure to itself.
Optimal Transport-based Anomaly Detection. Previous works (Alaoui-Belghiti et al., 2019, 2020) have considered OT for AD. These works proposed a distance-based detection mechanism, in which isolated samples are considered anomalies. OT contributes to this setting, by defining a rich metric between samples. Especially, these works considered AD in time series data, and OT is used to compute distances between those in the frequency domain, under a Chebyshev ground cost. In comparison with these methods, ours is notably general purpose, that is, we do not assume data to be time series. Instead of using a Chebyshev ground cost, we model the AD problem with a repulsive cost.
3 Proposed Method
3.1 Optimal Transport
OT is a field of mathematics, concerned with the displacement of mass between a source measure, and a target measure, at least effort. In the following, we cover the principles of OT in continuous and discrete settings. We refer readers to Peyré and Cuturi (2020) for a computational exposition of the main concepts, and Montesuma et al. (2024a) for applications in machine learning. In the following, we are particularly interested in the formulation by Kantorovich (1942), which is defined as,
Definition 3.1.
[…]
Equation 1 defines the transportation problem as an infinite dimensional linear program on the variable $\gamma$, called transport plan. In our case, instead of having access to a closed-form $P$, one has samples $\{\mathbf{x}_{i}^{(P)}\}_{i=1}^{n}$, each $\mathbf{x}_{i}^{(P)}\sim P$ with probability $p_{i}$. In such cases, $P$ may be approximated with an empirical measure,
$\hat{P}(\mathbf{x})=\sum_{i=1}^{n}p_{i}\delta(\mathbf{x}-\mathbf{x}_{i}^{(P)}).$ (2)
Since the $p_{i}^{\prime}s$ represent the probability of sampling $\mathbf{x}_{i}^{(P)}$, $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$. Plugging back equation 2 into equation 1 leads to a finite linear program,
$\hat{\gamma}=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{q})}{\text{arg min}}\,\langle\gamma,\mathbf{C}\rangle_{F}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}\underbrace{c(\mathbf{x}_{i}^{(P)},\mathbf{x}_{j}^{(Q)})}_{C_{ij}},$ (3)
where the optimization variables are the coefficients $\gamma_{ij}$ of the transport plan, and $\gamma\in\Gamma(\mathbf{p},\mathbf{q})$ indicates that $\sum_{i}\gamma_{ij}=q_{j}$ and $\sum_{j}\gamma_{ij}=p_{i}$. As a linear program, the solution $\gamma$ is a sparse matrix with at most $n+m-1$ non-zero elements (Peyré and Cuturi, 2020). Solving it through the celebrated Simplex algorithm (Dantzig, 1983), which has computational complexity $\mathcal{O}(n^{3}\log n)$ and storage complexity $\mathcal{O}(n^{2})$ (i.e., storing each $\gamma_{ij}$).
A faster alternative was introduced by Cuturi (2013), who shown that adding an entropic regularization to equation 3 leads to a problem that can be solved through Sinkhorn’s algorithm Sinkhorn (1967). From a continuous perspective, this is equivalent to penalizing Kullback-Leibler (KL) divergence between $\gamma$, and the trivial coupling $P\otimes Q=P(x)Q(y)$. This regularization term is related to the entropy of $\gamma$ as discussed in (Peyré and Cuturi, 2020, Chapter 4), hence this problem is called entropic OT.
Definition 3.2.
(Entropic Optimal Transport) Under the same conditions of Definition 3.1, let $\epsilon\geq 0$ be an entropic penalty. The entropic OT problem is given by,
[…]
where $\text{KL}(\gamma|\xi)=\int_{\mathcal{X}\times\mathcal{X}}\log\biggr{(}\dfrac{d\gamma}{d\xi}(x,y)\biggr{)}d\gamma(x,y)+\int_{\mathcal{X}\times\mathcal{X}}(d\xi(x,y)-d\gamma(x,y))$ is the KL divergence between measures $\gamma$ and $\xi$.
We can obtain an equivalent discrete formulation by plugging back equation 2 into 4, which leads to,
$\hat{\gamma}_{\epsilon}=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{q})}{\text{arg min}}\,\langle\gamma,\mathbf{C}\rangle_{F}-\epsilon H(\gamma)=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}C_{ij}+\epsilon\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}(\log\gamma_{ij}-1),$ (5)
where $H(\gamma)$ denotes the entropy of the transportation plan. Since equation 5 relies on the Sinkhorn (1967) algorithm rather than linear programming, it has $\mathcal{O}(Ln^{2})$ computational complexity, where $L$ is the number of iterations. In general, the KL and entropic terms in equations 4 and 5 has a smoothing effect over the joint measure $\gamma_{\epsilon}$. As a result, $\hat{\gamma}_{\epsilon}$ has more non-zero elements than $\hat{\gamma}$.
In the next section, we explore a new transportation with a single probability measure. This problem is understood as the transportation of $P$ to itself, when the samples form $P$ are forced to send their mass outside their immediate neighborhood.
3.2 Mass Repulsing Optimal Transport
In this section, we propose a new OT problem called MROT. This problem is inspired by the Kantorovich formulation, described in section 3.1. However, instead of considering two different probability measures $P,Q$, it considers the transportation of $P$ to itself. Due the properties of OT, if we consider the OT plan $\gamma^{\star}={\text{OT}}(P,P)$, it is supported on the set $\{(x,x),x\in\mathcal{X}\}$, i.e., each point keeps its own mass (Santambrogio, 2015). This motivates our new problem, in which we force points to repell its mass. We do so, by engineering the ground-cost $c:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$. For pairs $x\in\mathcal{X}$, $y\in\mathcal{X}$,
$\tilde{c}_{k}(x,y)=\begin{cases}c(x,y)&\text{ if }y\not\in\mathcal{N}_{k}(x),\\ \underset{z\in\mathcal{N}_{k}(x)}{\text{sup}}\,c(x,z)&\text{ otherwise}.\end{cases}$ (6)
this principle is different from repulsive costs (Di Marino et al., 2017), which are designed to model the interaction between particles in multi-marginal OT. Indeed, in our work we are designing a transportation problem from a probability measure to itself. While repulsive costs incentive transportation towards distant points in space, our mass repulsing cost induces transportation just outside an exclusion zone.
While the mass repulsing principle can be applied to any ground-cost $c$, let us focus on the metric case, i.e., when $c(x,y)=d(x,y)^{2}$, where $(\mathcal{X},d)$ is a metric space. The ground-cost we are proposing is essentially defining an exclusion zone, i.e., the $k-$vicinity $\mathcal{N}_{k}(x)$ centered at $x$, where $x$ cannot send its mass. Of course, since $x\in\mathcal{N}_{k}(x)$, it is forced to send its mass elsewhere – thus, we call this idea mass repulsive OT.
Our main hypothesis for anomaly detection is that anomalous points lie in low density regions of $P$. On the one hand, If these points are forced to send its mass outside its vicinity, it will be forced to send it to high density regions of $P$ – otherwise, mass conservation in OT would not hold. On the other hand, if the exclusion zone is smaller than high density regions of $P$, points on these regions will be sent close-by. As a result, anomalous points will have a higher transportation cost, given by $\gamma(x,y)\tilde{c}_{k}(x,y)$, than normal points. We thus consider the following OT problem,
[…]
which, like the continuous entropic OT problem in definition 3.2, also admits a discrete version when $P$ is approximated empirically through equation 2,
$\gamma^{\star}_{k,\epsilon}$ $=\text{MROT}_{k,\epsilon}(\hat{P})=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{p})}{\text{min}}\,\langle\gamma,\tilde{\mathbf{C}_{k}}\rangle_{F}-\epsilon H(\gamma).$ (8)
To ground-up intuition about our proposed problem, we show in Figure 1 a comparison between the commonly used squared Euclidean cost, our engineered cost, and a repulsive cost associated with the Coulomb interaction (Di Marino et al., 2017). In traditional OT, sending mass to far away regions of space is costly. As a result, points are encouraged to keep their own mass as close as possible. Consequently, OT from $P$ to itself is the trivial plan $\gamma=\text{Id}$.
Figure 1: Comparison between different ground costs for $x_{0}=(0,0)$. From left to right: Squared euclidean cost $c(x,x_{0})=\lVert x-x_{0}\rVert_{2}^{2}$, Engineered cost (c.f., equation 6) and repulsive cost.
Both our engineered cost $\tilde{c}_{k}$ and the repulsive cost $(1-c(x,x_{0}))^{-1}$ assign high costs to points near $x_{0}$. In particular, within the nearest-neighbor radius $\mathcal{N}{k}(x_{0})$, our engineered cost $\tilde{c}{k}(x,x_{0})$ reaches its maximum. However, there is an important difference between these costs. Our engineered cost promotes transportation to the immediate region outside $\mathcal{N}_{k}$, whereas the repulsive cost drives samples as far as possible from $x_{0}$. As we discussed previously, this feature of the ground-cost allows us to build an anomaly score.
3.3 Building and Generalizing an Anomaly Score
Through MROT, we want to build a score for samples, based on how anomalous they are. Assuming some samples lie in a low density region of space, our MROT problem forces those samples to send their mass to distant parts of the feature space. In contrast, normal samples can send their mass to the immediate neighborhood outside the exclusion zone. As a result, we can sort out normal from anomalous samples using the transportation effort,
[…]
where $\gamma(y|x)$ corresponds to the conditional probability, calculated through the joint $\gamma(x,y)$, given $x$. For empirical measures, this quantity can be calculated as follows,
$\mathcal{T}_{i}=\sum_{j=1}^{n}\dfrac{\gamma_{ij}^{\star}}{p_{i}}\tilde{c}_{k}(x_{i}^{(P)},x_{j}^{(P)}),$ (9)
where $p_{i}$ is the importance of the $i-$th sample (e.g, $p_{i}=n^{-1}$ for uniform importance). Interestingly, equation 9 is similar to the barycentric map, widely used in domain adaptation Courty et al. (2016); Montesuma et al. (2023). $\mathcal{T}$ has 2 shortcomings as an anomaly score. First, it is hardly interpretable, as its range depends on the choice of ground-cost $c$. Second, it is only defined in the support of $\hat{P}$. We offer a solution to both of these problems.
Concerning interpretability, we propose to transform it using the Cumulative Distribution Function (CDF) of its values. Let $P_{\mathcal{T}}$ be the probability measure associated with $\mathcal{T}(x)$. The CDF is simply $F_{\mathcal{T}}(t)=P_{\mathcal{T}}((-\infty,t))$. Naturally, since $P_{\mathcal{T}}$ is not available, it may be approximated from samples $\{\mathcal{T}_{i}\}_{i=1}^{n}$, obtained through equation 9. We do so through, Kernel Density Estimation (KDE),
$\hat{k}(t)$ $=\dfrac{1}{n\sigma}\sum_{i=1}^{n}\phi\biggr{(}\dfrac{t-\mathcal{T}_{i}}{\sigma}\biggr{)},$ (10)
where $K$ is a kernel function (e.g., the Gaussian kernel $\phi(x)=\exp(-\nicefrac{{x^{2}}}{{2}})$), and $\sigma$ is the bandwidth parameter, controlling the smoothness of $\hat{k}$ and determined through Scott’s rule (Scott, 1979). Equation 10 provide an approximation for the density of transportation efforts $\{\mathcal{T}_{i}\}_{i=1}^{n}$. The CDF can be easily calculated from this estimate, via,
[…]
The CDF is appealing, as it is a monotonic function over transportation efforts $t\in\mathbb{R}$, and it takes values on $[0,1]$, both of which are desirable properties for an anomaly score.
The issue on how to extrapolate the anomaly score for new samples remain. For instance, even if we use the CDF values as anomaly scores, one needs to recalculate $t=\hat{\mathcal{T}}(\mathbf{x})$ for a new sample, before evaluating $\hat{H}(t)$ or $\hat{K}(t)$. As we previously discussed, this is challenging, as $\hat{\mathcal{T}}$ is only defined on the support of $\hat{P}$. A naive approach would be to append $\mathbf{x}$ to the set $\{\mathbf{x}_{i}^{(P)}\}_{i=1}^{n}$ and solve a MROT problem again. Naturally, this is not feasible, as solving an OT problem for each new sample is computationally expensive.
In this paper we put forth the more efficient idea of modeling the relationship $\mathbf{x}\mapsto\hat{K}(\hat{\mathcal{T}}(\mathbf{x}))$ from the samples we are given, that is, we create a labeled dataset $\{\mathbf{x}_{i}^{(P)},t_{i}\}_{i=1}^{n}$, where $t_{i}=\hat{\mathcal{T}}(\mathbf{x}_{i}^{(P)})$. Our anomaly score comes, then, through a function $\psi(\mathbf{x}_{i}^{(P)})$ fit to the labeled dataset through regression. Fitting $\psi$ can be done with standard regression tools, such as Ordinary Least Squares (OLS), Support Vector Regression (SVR) (Smola and Schölkopf, 2004), nearest neighbors, or gradient boosting (Friedman, 2002). In general, one can expect the relationship between $\mathbf{x}_{i}^{(P)}$ and $\hat{K}(\hat{\mathcal{T}}(\mathbf{x}_{i}^{(P)}))$ to be non-linear, hence, it is generally necessary to use a non-linear regression model. We show a summary of our strategy in Algorithm 1.
1 function mrot($\mathbf{X}^{(P_{S})},\epsilon,k$)
2       $\gamma_{k,\epsilon}^{\star}=\text{MROT}_{k,\epsilon}(\hat{P})$;
3       $t_{i}\leftarrow\sum_{j=1}^{n}(\nicefrac{{(\gamma_{k,\epsilon}^{\star})_{ij}}}{{p_{i}}})\tilde{c}_{k}(x_{i}^{(P)},x_{j}^{(P)})$;
4       $\hat{k}\leftarrow\text{KDE}(\{t_{i}\}_{i=1}^{n})$;
5       $\psi\leftarrow\text{Regression}(\{x_{i}^{(P)},\hat{K}(t_{i})\}_{i=1}^{n})$;
6       return $\psi$;
7
Algorithm 1 Mass Repulsive Optimal Transport.
4 Experiments
We divide our experiments in 4 parts. Section 4.1 shows a toy example, going through all the steps in our algorithm. Section 4.2 shows our results on AdBench (Han et al., 2022). Section 4.3 shows our experiments in fault detection on the Tennessee Eastman Process (Montesuma et al., 2024b; Reinartz et al., 2021). Finally, Section 4.4 explores the robustness of our methods to various hyper-parameters and design choices.
4.1 An introductory toy example
Before diving into comparing our method with prior art, we give an introductory example that illustrates how we create anomaly scores out of samples. In this example, we sample normal examples $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},0.25\mathbf{I}_{2})$, and anomalous samples $\mathbf{y}_{j}\sim\mathcal{N}([-3,-3],0.01\mathbf{I}_{d})$, where $\mathbf{I}_{2}$ is a $2\times 2$ identity matrix. The dataset for this toy example consists of the concatenation $\{\mathbf{x}_{1},\cdots,\mathbf{x}_{n},\mathbf{y}_{1},\cdots,\mathbf{y}_{m}\}$, where $n=500$ and $m=25$, which means that roughly $5\%$ of the total number of samples are anomalies. In the following, we compare our engineered cost with a regularized Coulomb interaction, $C_{ij}=(1+\lVert\mathbf{x}_{i}^{(P)}-\mathbf{x}_{j}^{(P)}\rVert_{2})^{-1}$. This cost is shown alongside our proposed engineered cost in Figure 2. Note that, in Figures 2 (b - d) the lower-right corner of the matrices correspond to anomalies.
(a) Toy example.
(b) Euclidean cost.
(c) Engineered cost.
(d) Coulomb cost.
Figure 2: In (a), we show the samples for our toy example. In (b), (c) and (d), we show the ground-cost $C_{ij}$ between samples in (a), using the Euclidean distance, our engineered ground-cost (equation 6), and the regularized Coulomb interaction cost.
Figure 2 clearly illustrates that directly using the Euclidean cost results in a trivial OT solution, because the diagonal entries of $C_{ij}$ are zero. This is not the case for our engineered cost, and the regularized Coulomb interaction, shown in Figures 2 (c) and (d). To stress this idea, we show in Figure 3 the transportation plans acquired by MROT and using the regularized Coulomb cost. While both strategies achieve similar transportation matrices, in the second case the anomalies send and receive less mass than normal samples. This is somewhat expected, as the anomalous samples are clustered together in a tight region of $\mathbb{R}^{2}$, resulting in a high cost. As a result, these samples are encouraged to send their mass elsewhere.
(a) MROT.
(b) OT with the regularized Coulomb cost.
Figure 3: In (a), we show the MROT transportation plan (left side) and strategy (right side). Likewise, in (b) we show the transportation plan and strategy for the regularized Coulomb cost. While both transportation strategies achieve similar results, in (b) the anomalies (lower right corner) sends and receives much less mass than other normal samples.
Here comes an important distinction between our approach and simply using repulsive costs in OT. Considering our last remark about the regularized Coulomb interaction, samples send their mass to distant regions of space, thus reducing $C_{ij}$. As a result, anomalies will incur in a smaller transportation cost than normal samples. Although counterintuitive, one can still construct a detection rule out of this idea, as we show below.
In Figure 4, we present the anomaly scores derived from our MROT strategy and the regularized Coulomb cost. These scores are computed by first determining the transportation effort (c.f., equation 9) for each sample $x_{i}^{(P)}$, then estimating the density of these efforts, and finally transforming the density into a $[0,1]$ score via its CDF, as outlined in section 3.3. As shown in Figure 4, the MROT approach assigns higher anomaly scores to anomalous samples compared to normal ones, simplifying the process of setting a threshold for anomaly detection. In contrast, the regularized Coulomb cost exhibits the opposite behavior. Indeed, anomalous samples send their mass to distant parts of space (i.e., to normal samples), which, due the nature of the Coulomb cost, lead to a smaller transportation effort. Nevertheless, as discussed previously, it is still possible to establish a detection rule, albeit being counterintuitive.
(a) MROT anomaly score.
(b) regularized Coulomb cost anomaly score.
Figure 4: Anomaly score comparison between the MROT (ours, a) and OT with repulsive costs (b). Scores are computed by calculating the transportation effort (Eq. 9) for each sample $x_{i}^{(P)}$, estimating their density, and normalizing to the $[0,1]$ range via the CDF (c.f., section 3.3). Overall, our strategy assigns higher scores (close to $1.0$) to anomalous samples, while assigning smaller scores to normal samples.
As we discussed throughout section 3.3, the scores in Figure 4 are only defined in the support of $\hat{P}$. We then explore the regression of this score through regression, which can be done with a variety of standard regression algorithms. In Figure 5 we show the results for eXtreme Gradient Boosting (XGBoost) (Chen and Guestrin, 2016) and SVR (Smola and Schölkopf, 2004), which both define the anomaly score over the whole ambient space $\mathbb{R}^{2}$. As an important remark, the relationship between $x_{i}^{(P)}$ and its score is likely non-linear. Indeed, this idea is evidenced in Figure 5. As a result, one needs a non-linear regression algorithm.
(a) MROT regressed score.
(b) regularized Coulomb cost regressed score.
Figure 5: Regression of the anomaly score for MROT (ours, a) and OT with repulsive costs (b).
4.2 Comparison on AdBench
AdBench (Han et al., 2022) is a benchmark in AD with 57 different kinds of datasets, grouped into 47, 5 and 5 real-world, vision and Natural Language Processing (NLP) datasets. In our experiments, we focus on real-world and natural-language datasets. We compare, in total, 17 methods, grouped into classical, diffusion-based and our proposed methods based on OT. For classic methods, we consider Isolation Forest (IsoF) (Liu et al., 2008), One Class SVM (OCSVM) (Schölkopf et al., 1999), $k-$NN, LOF (Breunig et al., 2000), CBLOF (He et al., 2003), Empirical Cumulative Distribution Functions (ECOD) (Li et al., 2022), Copula-Based OD (COPOD) (Li et al., 2020), Lightweight On-line Detector of Anomalies (LODA) (Pevnỳ, 2016), Feature Bagging (Lazarevic and Kumar, 2005), and Histogram-based Outlier score (HBOS) (Goldstein and Dengel, 2012). For diffusion based, we consider the 4 variants of Diffusion Time Estimation (DTE) (Livernoche et al., 2024). Finally, we consider OT with repulsive costs Di Marino et al. (2017), and our proposed MROT.
(a) Real-world.
(b) NLP
Figure 6: AdBench result summary. AUC-ROC per dataset is available in the appendix. We compare, in total, 17 algorithms over 47 real-world datasets (a) and 5 NLP datasets (b). For real-world datasets, our MROT has state-of-the-art performance, especially when compared with recently proposed flow-based models. For higher dimensional data, our method
We show our summarized results in Figures 6 (a) and (b), for real-world and NLP datasets, respectively. First, we note that MROT and OT with repulsive costs have superior performance with respect other methods on real-world datasets. This result highlights the usefulness of OT theory in the analysis of probability distributions. Furthermore, on average, our MROT has better performance than OT (77.36% versus 75.97% ROC-AUC), proving the effectiveness of our cost engineering strategy. We refer readers to our appendix for detailed results per dataset.
Even though our method has superior performance on real-world datasets, the NLP benchmarks reveal limitations of OT based AD. A first limitation of using MROT comes from the scalability of OT. As a linear program, it has at least $\mathcal{O}(n^{2})$ storage, and $\mathcal{O}(n^{3}\log n)$ computational complexity, where $n$ is the number of samples. In our experiments, we limited the number of samples to $n=20,000$, by down-sampling bigger datasets. As reported in Figure 6 (a), and our detailed results are given in the Appendix, this process does not affect performance. Furthermore, from Figure 6 (b), we see that our method struggles in high dimensional anomaly detection, such as those in the NLP datasets of Han et al. (2022). This limitation stems from the use of OT in high-dimensions (Montesuma et al., 2025, Section 8.1).
4.3 Tennessee Eastman Process
In this section, we focus in comparing anomaly detection methods for fault detection. To that end, we use the Tennessee Eastmann (TE) process benchmark (Downs and Vogel, 1993), especially the simulations of Reinartz et al. (2021), pre-processed by Montesuma et al. (2024b). This benchmark is composed by simulations of a large-scale chemical plant. Besides the normal state, there are 28 faults associated with different parts of the plant. An interesting feature of this benchmark is that there are different modes of operation. Ideally, an anomaly detection algorithm trained in one mode should generalize to other modes, but this is not guaranteed due to distributional shift. Such challenge is the subject of domain adaptation (Montesuma et al., 2023, 2024b, 2024c, 2025), which is beyond the scope of this work. We refer readers to the aforementioned papers for further details.
From the original simulations in Reinartz et al. (2021), we consider $100$ simulations for the normal state. These simulations last for $600$ hours, with a time-step of $1$ hour, and concern $34$ sensors measuring different physical and chemical properties. On top of these $100$ simulations, we take $1$ faulty simulation for each kind of fault, leading to a total of $128$ simulations. Based on this set of simulations, we extract windows of $20$ hours from the original signals. Each window is then considered a sample for anomaly detection. On each window, we compute the mean, and standard deviation of each variable, leading to vectors $\mu,\sigma\in\mathbb{R}^{34}$. We use the concatenation of these vectors as features for anomaly detection. These steps lead to $6$ datasets with $3840$ samples, $840$ ($21.875\%$) of them being anomalous, one for each mode of operation. We thus divide our discussion into two parts: fault detection, and cross-domain fault detection.
(a) Mode 1.
(b) Mode 2.
(c) Mode 3.
(d) Mode 4.
(e) Mode 5.
(f) Mode 6.
Figure 7: t-SNE embeddings of the Tennessee Eastman process data per mode. In blue, we show the normal samples, whereas we show anomalous samples in shades of reds, corresponding to the actual fault category they correspond. While most anomalous samples do cluster in a region outside the non-faulty cluster, some faulty samples do not.
In Figure 7, we visualize the Tennessee Eastman data using t-SNE. We scatter the embeddings of each mode’s data, showing that most faulty samples cluster outside the normal data cluster. However, some faulty samples are close to normal ones. This phenomenon is expected, as the effect of faults evolves over time. As a result, windows taken in early stages of simulation resemble those of normal samples.
Next, we benchmark the AD performance of algorithms, and their capability to generalize to unseen that within the same mode of operation. In this experiment, we downsample the number of anomalous samples per fault category to $\{5,10,\cdots,30\}$. This results in a percentage of $\{4.45\%,8.53\%,12.28\%,15.73\%,18.92\%,21.87\%\}$ of anomalous samples. In Figure 8, we report our aggregated results over all percentage of anomalies. We refer readers to our appendix for results per percentage.
The main idea of this methodology is evaluating how different algorithms perform, under a variable percentage of faults. Our results are shown in Figure 7 (b). Among the tested methods, MROT has a better performance and remains stable throughout the range of percentage of anomalies.
(a) Mode 1.
(b) Mode 2.
(c) Mode 3.
(d) Mode 4.
(e) Mode 5.
(f) Mode 6.
Figure 8: Aggregated anomaly detection on the Tennessee Eastman data per mode of operation. First, MROT outperforms OT with repulsive costs over all modes. Second, MROT and DTE-C have state-of-the-art performance, superior to previously proposed methods.
4.4 Robustness
(a) MROT.
(b) ROT.
Figure 9: Hyper-parameter sensitivity. In (a), we show the performance of MROT for a variable number of nearest neighbors $k$, and entropic regularization $\epsilon$. Overall, our method is robust to the choices of these hyper-parameters, but using a lower $\epsilon$ is generally better. In contrast, we show in (b) the performance of OT-based AD with the regularized Coulomb cost, for which using a higher entropic penalty $\epsilon$ improves performance.
Our method has two hyper-parameters, the entropic regularization penalty $\epsilon$, and the number of nearest neighbors $k$. In our experiments, we evaluated our methods over the values $\epsilon\in\{0,10^{-2},10^{-1},10^{0}\}$, where $\epsilon=0$ implies using exact OT, i.e., linear programming. For MROT, we use $k\in\{5,10,20,\cdots,50\}$. Note that, while $\epsilon$ is linked to the transportation plan, $k$ is linked to the ground-cost. We summarize our results in Figure 9, where we present KDEs over the AUC-ROC scores of each dataset, for each combination of hyper-parameters.
From Figure 9, we note that both MROT is robust to the choice of entropic regularization and number of nearest neighbors. As a general guideline, it is better to limit the number of nearest neighbors, as anomalous examples are likely rare. As a consequence, using a high value for $k$ may lead to the inclusion of normal points in the neighborhood of anomalous ones.
5 Conclusion
In this paper, we introduced a novel, general purpose anomaly detection algorithm based on optimal transport theory. Our method works under the assumption that the local neighborhood of anomalous samples is likely more irregular than that of normal samples. Based on this idea, we engineer the ground-cost in optimal transport, to encourage samples to send their mass just outside an exclusion zone, defined through its $k-$nearest neighbors. While this idea bears some similarity to optimal transport with repulsive costs (Di Marino et al., 2017), the defined ground-cost is not repulsive, and, as we show in our experiments, it leads to better anomaly detectors. We thoroughly experiment on the AdBench (Han et al., 2022) benchmark, and the Tennessee Eastman process (Downs and Vogel, 1993; Reinartz et al., 2021; Montesuma et al., 2024b), showing that our method outperforms previously proposed methods. However, as we highlight in section 4.2, our method inherits the limitations of classic optimal transport (Montesuma et al., 2024a). Especially, it needs $\mathcal{O}(n^{2})$ computational storage, has $\mathcal{O}(n^{3}\log n)$ computational complexity and is difficult to estimate in high dimensions. Nonetheless, as our experiments demonstrate, our method is practical and performative in real-world anomaly detection scenarios.
Appendix A Additional Details on Experiments
A.1 Additional results on AdBench
Table 1: Comparison of AUC-ROC results on ADBench.
Dataset IsoF OCSVM k-NN PCA LOF CBLOF ECOD COPOD LODA FeatureBagging HBOS OT MROT DTE-IG DTE-NP DDPM DTE-C cover 91.48 $\pm$ 1.50 65.64 $\pm$ 0.25 88.61 $\pm$ 0.39 83.12 $\pm$ 0.34 56.75 $\pm$ 1.89 92.24 $\pm$ 0.17 92.02 $\pm$ 0.49 88.20 $\pm$ 0.32 92.18 $\pm$ 3.85 57.14 $\pm$ 2.15 70.68 $\pm$ 1.16 92.61 $\pm$ 0.28 88.01 $\pm$ 1.01 82.76 $\pm$ 1.85 81.76 $\pm$ 1.75 80.76 $\pm$ 1.65 83.76 $\pm$ 1.95 donors 78.24 $\pm$ 1.15 70.11 $\pm$ 0.14 71.91 $\pm$ 0.20 58.16 $\pm$ 0.19 62.89 $\pm$ 1.35 80.77 $\pm$ 0.71 88.87 $\pm$ 0.11 81.53 $\pm$ 0.27 56.62 $\pm$ 38.02 69.06 $\pm$ 1.74 74.31 $\pm$ 0.63 84.85 $\pm$ 0.21 84.51 $\pm$ 0.19 82.63 $\pm$ 2.08 81.63 $\pm$ 1.98 80.63 $\pm$ 1.88 83.63 $\pm$ 2.18 fault 57.42 $\pm$ 3.35 53.42 $\pm$ 1.30 70.27 $\pm$ 2.38 48.37 $\pm$ 2.12 57.88 $\pm$ 1.35 66.50 $\pm$ 2.51 45.35 $\pm$ 1.77 45.49 $\pm$ 0.16 47.78 $\pm$ 2.34 59.10 $\pm$ 1.16 50.62 $\pm$ 6.66 54.29 $\pm$ 3.10 56.35 $\pm$ 7.46 58.22 $\pm$ 1.15 57.22 $\pm$ 1.05 56.22 $\pm$ 0.95 59.22 $\pm$ 1.25 fraud 94.47 $\pm$ 1.36 94.83 $\pm$ 0.38 96.57 $\pm$ 0.80 90.31 $\pm$ 0.36 54.77 $\pm$ 7.30 95.39 $\pm$ 0.81 94.87 $\pm$ 0.81 94.29 $\pm$ 1.42 85.55 $\pm$ 5.41 61.59 $\pm$ 7.15 94.51 $\pm$ 1.22 95.08 $\pm$ 1.21 94.48 $\pm$ 1.37 94.37 $\pm$ 1.57 93.37 $\pm$ 1.47 92.37 $\pm$ 1.37 95.37 $\pm$ 1.67 glass 80.49 $\pm$ 12.70 45.44 $\pm$ 4.93 76.34 $\pm$ 9.85 55.13 $\pm$ 5.13 61.76 $\pm$ 12.42 85.50 $\pm$ 4.18 75.37 $\pm$ 13.70 75.95 $\pm$ 1.51 62.43 $\pm$ 12.75 65.86 $\pm$ 12.47 82.02 $\pm$ 1.77 80.24 $\pm$ 10.54 78.29 $\pm$ 10.58 58.04 $\pm$ 9.31 57.04 $\pm$ 9.21 56.04 $\pm$ 9.11 59.04 $\pm$ 9.41 Hepatitis 62.05 $\pm$ 13.25 48.22 $\pm$ 2.80 46.67 $\pm$ 13.62 51.67 $\pm$ 6.77 33.75 $\pm$ 1.07 99.61 $\pm$ 0.03 69.23 $\pm$ 6.78 99.07 $\pm$ 0.26 5.95 $\pm$ 9.35 28.82 $\pm$ 1.45 99.11 $\pm$ 0.16 84.62 $\pm$ 11.03 81.54 $\pm$ 14.66 99.96 $\pm$ 0.36 99.86 $\pm$ 0.26 99.76 $\pm$ 0.16 100.06 $\pm$ 0.46 http 99.96 $\pm$ 0.02 99.53 $\pm$ 0.01 9.41 $\pm$ 0.55 95.15 $\pm$ 0.06 54.86 $\pm$ 1.33 54.77 $\pm$ 4.28 97.85 $\pm$ 0.02 42.17 $\pm$ 0.12 38.23 $\pm$ 3.78 54.03 $\pm$ 1.15 57.50 $\pm$ 0.58 99.46 $\pm$ 0.06 99.31 $\pm$ 0.31 51.62 $\pm$ 0.79 50.62 $\pm$ 0.69 49.62 $\pm$ 0.59 52.62 $\pm$ 0.89 InternetAds 70.93 $\pm$ 2.22 69.77 $\pm$ 1.82 67.05 $\pm$ 2.71 58.00 $\pm$ 1.40 87.81 $\pm$ 0.48 76.32 $\pm$ 1.80 67.74 $\pm$ 2.69 56.02 $\pm$ 0.07 53.71 $\pm$ 4.29 88.58 $\pm$ 0.51 58.85 $\pm$ 0.59 65.48 $\pm$ 3.77 68.37 $\pm$ 5.97 86.73 $\pm$ 1.64 85.73 $\pm$ 1.54 84.73 $\pm$ 1.44 87.73 $\pm$ 1.74 Ionosphere 83.48 $\pm$ 1.91 82.61 $\pm$ 2.58 85.36 $\pm$ 2.05 67.22 $\pm$ 3.64 67.84 $\pm$ 0.38 72.53 $\pm$ 0.10 75.37 $\pm$ 6.49 68.10 $\pm$ 0.04 65.50 $\pm$ 1.41 69.96 $\pm$ 0.64 70.93 $\pm$ 0.55 82.90 $\pm$ 2.11 81.91 $\pm$ 3.27 78.26 $\pm$ 1.68 77.26 $\pm$ 1.58 76.26 $\pm$ 1.48 79.26 $\pm$ 1.78 landsat 47.52 $\pm$ 1.57 41.31 $\pm$ 1.28 61.83 $\pm$ 1.88 46.74 $\pm$ 0.55 70.15 $\pm$ 1.87 79.50 $\pm$ 1.84 37.35 $\pm$ 1.10 90.54 $\pm$ 0.03 86.69 $\pm$ 2.45 72.61 $\pm$ 0.51 83.77 $\pm$ 0.90 43.51 $\pm$ 1.58 67.72 $\pm$ 1.78 76.94 $\pm$ 2.34 75.94 $\pm$ 2.24 74.94 $\pm$ 2.14 77.94 $\pm$ 2.44 ALOI 53.27 $\pm$ 1.93 52.04 $\pm$ 0.40 60.38 $\pm$ 1.59 51.99 $\pm$ 0.36 65.77 $\pm$ 1.94 84.26 $\pm$ 1.11 52.10 $\pm$ 1.87 50.00 $\pm$ 0.00 56.40 $\pm$ 11.07 66.44 $\pm$ 1.42 57.36 $\pm$ 0.38 54.19 $\pm$ 2.11 53.69 $\pm$ 1.68 83.60 $\pm$ 1.40 82.60 $\pm$ 1.30 81.60 $\pm$ 1.20 84.60 $\pm$ 1.50 letter 63.35 $\pm$ 6.64 50.91 $\pm$ 1.92 82.54 $\pm$ 1.27 51.31 $\pm$ 0.37 58.07 $\pm$ 8.89 100.00 $\pm$ 0.00 58.30 $\pm$ 2.73 94.81 $\pm$ 0.48 99.30 $\pm$ 0.77 57.54 $\pm$ 10.67 100.00 $\pm$ 0.00 53.11 $\pm$ 4.87 70.61 $\pm$ 5.33 99.95 $\pm$ 0.23 99.95 $\pm$ 0.13 99.95 $\pm$ 0.03 99.95 $\pm$ 0.33 Lymphography 100.00 $\pm$ 0.00 99.37 $\pm$ 0.32 98.62 $\pm$ 3.08 94.56 $\pm$ 1.30 53.82 $\pm$ 5.25 78.48 $\pm$ 0.95 98.62 $\pm$ 1.89 50.00 $\pm$ 0.00 49.29 $\pm$ 9.35 53.93 $\pm$ 5.59 86.81 $\pm$ 0.25 91.03 $\pm$ 13.04 99.31 $\pm$ 1.54 42.15 $\pm$ 3.91 41.15 $\pm$ 3.81 40.15 $\pm$ 3.71 43.15 $\pm$ 4.01 magic.gamma 71.93 $\pm$ 0.58 67.25 $\pm$ 0.35 79.89 $\pm$ 0.92 60.08 $\pm$ 0.27 53.42 $\pm$ 4.61 86.38 $\pm$ 7.22 64.40 $\pm$ 0.97 90.60 $\pm$ 0.43 89.51 $\pm$ 0.85 51.82 $\pm$ 5.22 92.47 $\pm$ 0.38 69.54 $\pm$ 0.99 71.54 $\pm$ 0.89 72.03 $\pm$ 5.63 71.03 $\pm$ 5.53 70.03 $\pm$ 5.43 73.03 $\pm$ 5.73 mammography 84.95 $\pm$ 4.51 87.00 $\pm$ 0.53 86.37 $\pm$ 1.72 75.40 $\pm$ 0.49 70.95 $\pm$ 1.05 67.57 $\pm$ 0.98 90.33 $\pm$ 2.46 77.67 $\pm$ 0.17 45.33 $\pm$ 12.81 78.77 $\pm$ 2.68 60.84 $\pm$ 2.38 88.46 $\pm$ 2.27 88.24 $\pm$ 1.66 83.37 $\pm$ 1.57 82.37 $\pm$ 1.47 81.37 $\pm$ 1.37 84.37 $\pm$ 1.67 mnist 80.43 $\pm$ 2.20 82.28 $\pm$ 0.71 84.81 $\pm$ 0.77 67.44 $\pm$ 1.85 54.98 $\pm$ 1.29 74.19 $\pm$ 3.23 74.84 $\pm$ 1.00 63.34 $\pm$ 0.04 61.39 $\pm$ 2.92 54.54 $\pm$ 1.26 76.18 $\pm$ 0.55 86.49 $\pm$ 1.28 88.16 $\pm$ 1.05 73.53 $\pm$ 0.72 72.53 $\pm$ 0.62 71.53 $\pm$ 0.52 74.53 $\pm$ 0.82 musk 99.97 $\pm$ 0.05 100.00 $\pm$ 0.00 90.02 $\pm$ 3.10 95.91 $\pm$ 0.36 53.91 $\pm$ 7.29 99.88 $\pm$ 0.01 95.26 $\pm$ 0.80 97.46 $\pm$ 0.04 98.14 $\pm$ 0.49 52.55 $\pm$ 7.36 97.60 $\pm$ 0.12 100.00 $\pm$ 0.00 99.97 $\pm$ 0.06 99.75 $\pm$ 0.25 99.65 $\pm$ 0.15 99.55 $\pm$ 0.05 99.85 $\pm$ 0.35 optdigits 73.76 $\pm$ 5.09 54.84 $\pm$ 2.28 39.46 $\pm$ 3.03 45.23 $\pm$ 0.36 52.55 $\pm$ 1.09 62.06 $\pm$ 3.96 63.04 $\pm$ 2.53 99.50 $\pm$ 0.10 38.87 $\pm$ 13.16 49.29 $\pm$ 3.40 98.63 $\pm$ 0.23 55.07 $\pm$ 2.92 55.21 $\pm$ 1.13 97.70 $\pm$ 2.48 97.60 $\pm$ 2.38 97.50 $\pm$ 2.28 97.80 $\pm$ 2.58 PageBlocks 90.72 $\pm$ 1.82 68.81 $\pm$ 0.72 58.50 $\pm$ 2.43 72.16 $\pm$ 1.98 54.97 $\pm$ 0.17 67.51 $\pm$ 3.80 90.75 $\pm$ 0.89 47.12 $\pm$ 0.19 44.16 $\pm$ 1.92 53.38 $\pm$ 0.29 58.84 $\pm$ 0.37 90.29 $\pm$ 0.52 90.27 $\pm$ 0.72 48.09 $\pm$ 2.31 47.09 $\pm$ 2.21 46.09 $\pm$ 2.11 49.09 $\pm$ 2.41 pendigits 94.74 $\pm$ 1.12 95.81 $\pm$ 0.17 82.88 $\pm$ 1.73 78.44 $\pm$ 0.66 89.92 $\pm$ 5.31 86.29 $\pm$ 4.76 92.74 $\pm$ 2.46 91.17 $\pm$ 1.63 81.88 $\pm$ 3.56 79.36 $\pm$ 5.06 80.91 $\pm$ 4.76 94.80 $\pm$ 1.02 88.74 $\pm$ 1.09 95.76 $\pm$ 1.50 95.66 $\pm$ 1.40 95.56 $\pm$ 1.30 95.86 $\pm$ 1.60 Pima 67.95 $\pm$ 1.39 66.33 $\pm$ 1.73 62.42 $\pm$ 2.36 54.35 $\pm$ 0.85 51.15 $\pm$ 0.65 47.11 $\pm$ 0.24 53.95 $\pm$ 5.03 48.89 $\pm$ 0.29 46.57 $\pm$ 1.85 50.85 $\pm$ 0.71 47.32 $\pm$ 0.19 72.81 $\pm$ 1.23 72.72 $\pm$ 0.95 48.57 $\pm$ 1.37 47.57 $\pm$ 1.27 46.57 $\pm$ 1.17 49.57 $\pm$ 1.47 annthyroid 81.13 $\pm$ 1.56 57.20 $\pm$ 0.67 69.45 $\pm$ 2.57 60.50 $\pm$ 1.09 65.69 $\pm$ 3.08 90.93 $\pm$ 1.07 79.06 $\pm$ 1.70 93.91 $\pm$ 0.26 81.89 $\pm$ 12.49 70.68 $\pm$ 1.69 94.84 $\pm$ 1.01 67.49 $\pm$ 0.77 67.56 $\pm$ 3.28 89.06 $\pm$ 1.70 88.06 $\pm$ 1.60 87.06 $\pm$ 1.50 90.06 $\pm$ 1.80 satellite 69.06 $\pm$ 1.91 66.24 $\pm$ 0.37 72.76 $\pm$ 0.85 64.68 $\pm$ 1.51 48.68 $\pm$ 3.40 46.34 $\pm$ 1.41 58.69 $\pm$ 1.30 26.28 $\pm$ 3.94 29.44 $\pm$ 5.11 47.32 $\pm$ 5.60 31.73 $\pm$ 3.46 65.36 $\pm$ 1.00 74.69 $\pm$ 1.12 58.34 $\pm$ 7.61 57.34 $\pm$ 7.51 56.34 $\pm$ 7.41 59.34 $\pm$ 7.71 satimage-2 99.25 $\pm$ 0.94 99.62 $\pm$ 0.14 99.67 $\pm$ 0.25 92.63 $\pm$ 1.10 76.42 $\pm$ 2.72 89.71 $\pm$ 0.72 95.73 $\pm$ 2.03 50.00 $\pm$ 0.00 51.50 $\pm$ 16.40 79.03 $\pm$ 3.04 74.04 $\pm$ 0.77 99.50 $\pm$ 0.47 99.85 $\pm$ 0.20 91.18 $\pm$ 0.82 90.18 $\pm$ 0.72 89.18 $\pm$ 0.62 92.18 $\pm$ 0.92 shuttle 99.72 $\pm$ 0.11 99.19 $\pm$ 0.03 79.43 $\pm$ 0.48 96.46 $\pm$ 0.07 93.18 $\pm$ 1.07 88.38 $\pm$ 0.50 99.29 $\pm$ 0.05 49.60 $\pm$ 0.44 70.52 $\pm$ 4.82 93.29 $\pm$ 1.73 67.91 $\pm$ 0.98 99.32 $\pm$ 0.30 99.06 $\pm$ 0.09 92.34 $\pm$ 2.96 91.34 $\pm$ 2.86 90.34 $\pm$ 2.76 93.34 $\pm$ 3.06 skin 64.79 $\pm$ 1.18 54.38 $\pm$ 0.33 71.22 $\pm$ 0.05 43.68 $\pm$ 0.08 32.97 $\pm$ 14.63 45.27 $\pm$ 33.40 48.73 $\pm$ 0.14 86.46 $\pm$ 4.66 82.17 $\pm$ 3.91 32.34 $\pm$ 5.54 90.72 $\pm$ 3.31 61.75 $\pm$ 0.24 76.09 $\pm$ 0.17 39.39 $\pm$ 8.57 38.39 $\pm$ 8.47 37.39 $\pm$ 8.37 40.39 $\pm$ 8.67 smtp 85.78 $\pm$ 4.20 75.50 $\pm$ 3.03 92.28 $\pm$ 4.06 82.41 $\pm$ 2.03 45.30 $\pm$ 2.12 46.06 $\pm$ 1.23 90.46 $\pm$ 7.47 38.03 $\pm$ 0.15 46.13 $\pm$ 4.68 46.45 $\pm$ 1.53 40.17 $\pm$ 0.97 89.58 $\pm$ 5.74 89.82 $\pm$ 8.77 48.31 $\pm$ 0.92 47.31 $\pm$ 0.82 46.31 $\pm$ 0.72 49.31 $\pm$ 1.02 SpamBase 57.66 $\pm$ 1.63 51.24 $\pm$ 1.00 72.70 $\pm$ 1.02 48.89 $\pm$ 0.95 44.61 $\pm$ 3.25 96.08 $\pm$ 0.93 64.61 $\pm$ 0.59 99.44 $\pm$ 0.16 96.97 $\pm$ 3.27 40.83 $\pm$ 2.60 98.44 $\pm$ 0.30 57.07 $\pm$ 1.58 56.16 $\pm$ 1.62 78.64 $\pm$ 4.42 77.64 $\pm$ 4.32 76.64 $\pm$ 4.22 79.64 $\pm$ 4.52 speech 54.70 $\pm$ 10.76 46.74 $\pm$ 0.88 46.13 $\pm$ 8.62 50.40 $\pm$ 0.98 61.44 $\pm$ 0.34 73.78 $\pm$ 0.30 45.31 $\pm$ 9.31 78.28 $\pm$ 0.04 49.27 $\pm$ 8.77 59.37 $\pm$ 4.19 76.81 $\pm$ 0.27 58.57 $\pm$ 10.63 61.04 $\pm$ 8.76 74.38 $\pm$ 0.97 73.38 $\pm$ 0.87 72.38 $\pm$ 0.77 75.38 $\pm$ 1.07 Stamps 90.43 $\pm$ 1.49 85.28 $\pm$ 2.83 89.89 $\pm$ 3.53 63.15 $\pm$ 9.14 55.12 $\pm$ 2.42 83.16 $\pm$ 1.77 88.06 $\pm$ 4.22 92.08 $\pm$ 0.30 85.59 $\pm$ 7.10 57.89 $\pm$ 2.77 83.94 $\pm$ 1.23 90.27 $\pm$ 1.28 89.14 $\pm$ 5.02 74.33 $\pm$ 6.13 73.33 $\pm$ 6.03 72.33 $\pm$ 5.93 75.33 $\pm$ 6.23 thyroid 97.83 $\pm$ 1.10 89.35 $\pm$ 0.57 95.53 $\pm$ 1.17 87.00 $\pm$ 3.16 43.21 $\pm$ 1.22 75.34 $\pm$ 1.76 97.28 $\pm$ 0.54 75.70 $\pm$ 0.62 59.97 $\pm$ 11.91 51.39 $\pm$ 2.70 75.41 $\pm$ 0.65 95.88 $\pm$ 0.66 95.96 $\pm$ 1.65 81.58 $\pm$ 2.08 80.58 $\pm$ 1.98 79.58 $\pm$ 1.88 82.58 $\pm$ 2.18 vertebral 36.43 $\pm$ 7.36 44.36 $\pm$ 7.51 22.62 $\pm$ 11.14 43.15 $\pm$ 2.60 56.20 $\pm$ 0.61 66.40 $\pm$ 0.18 44.05 $\pm$ 6.31 50.00 $\pm$ 0.00 45.44 $\pm$ 13.11 53.75 $\pm$ 0.31 61.09 $\pm$ 0.34 26.35 $\pm$ 5.33 23.17 $\pm$ 10.55 67.87 $\pm$ 0.34 66.87 $\pm$ 0.24 65.87 $\pm$ 0.14 68.87 $\pm$ 0.44 backdoor 73.47 $\pm$ 3.27 86.32 $\pm$ 0.14 73.93 $\pm$ 0.54 84.20 $\pm$ 1.07 76.66 $\pm$ 0.35 55.58 $\pm$ 0.20 84.87 $\pm$ 0.58 51.53 $\pm$ 0.01 49.52 $\pm$ 1.02 79.15 $\pm$ 0.56 53.11 $\pm$ 0.21 88.77 $\pm$ 0.56 88.80 $\pm$ 0.84 55.22 $\pm$ 0.45 54.22 $\pm$ 0.35 53.22 $\pm$ 0.25 56.22 $\pm$ 0.55 vowels 77.30 $\pm$ 2.40 77.17 $\pm$ 1.34 93.79 $\pm$ 1.71 59.48 $\pm$ 2.51 52.67 $\pm$ 2.09 56.09 $\pm$ 2.47 56.87 $\pm$ 4.12 66.42 $\pm$ 3.42 70.80 $\pm$ 13.23 53.79 $\pm$ 1.73 59.50 $\pm$ 1.09 63.52 $\pm$ 3.49 93.55 $\pm$ 3.05 59.86 $\pm$ 4.47 58.86 $\pm$ 4.37 57.86 $\pm$ 4.27 60.86 $\pm$ 4.57 Waveform 71.16 $\pm$ 3.99 66.83 $\pm$ 2.93 78.04 $\pm$ 1.81 52.53 $\pm$ 0.90 46.77 $\pm$ 8.48 63.47 $\pm$ 14.61 61.45 $\pm$ 4.70 80.74 $\pm$ 0.93 55.74 $\pm$ 15.73 46.94 $\pm$ 11.14 76.77 $\pm$ 1.57 74.26 $\pm$ 3.29 71.09 $\pm$ 3.44 48.12 $\pm$ 8.23 47.12 $\pm$ 8.13 46.12 $\pm$ 8.03 49.12 $\pm$ 8.33 WBC 99.77 $\pm$ 0.52 99.36 $\pm$ 0.29 100.00 $\pm$ 0.00 95.44 $\pm$ 1.53 58.73 $\pm$ 1.53 61.55 $\pm$ 0.12 100.00 $\pm$ 0.00 67.61 $\pm$ 0.07 54.06 $\pm$ 5.45 49.42 $\pm$ 4.74 69.56 $\pm$ 0.04 98.84 $\pm$ 0.82 99.07 $\pm$ 0.52 63.37 $\pm$ 0.23 62.37 $\pm$ 0.13 61.37 $\pm$ 0.03 64.37 $\pm$ 0.33 WDBC 97.64 $\pm$ 3.17 99.83 $\pm$ 0.02 99.86 $\pm$ 0.31 95.70 $\pm$ 1.57 86.38 $\pm$ 2.12 89.19 $\pm$ 1.14 96.53 $\pm$ 2.35 78.27 $\pm$ 1.27 78.84 $\pm$ 4.28 87.64 $\pm$ 1.63 54.44 $\pm$ 3.14 98.61 $\pm$ 1.47 98.75 $\pm$ 0.91 77.75 $\pm$ 2.61 76.75 $\pm$ 2.51 75.75 $\pm$ 2.41 78.75 $\pm$ 2.71 Wilt 44.25 $\pm$ 3.33 39.65 $\pm$ 2.14 68.71 $\pm$ 3.11 45.61 $\pm$ 0.49 63.63 $\pm$ 18.01 99.35 $\pm$ 0.74 38.47 $\pm$ 4.20 99.60 $\pm$ 0.09 90.04 $\pm$ 11.06 52.34 $\pm$ 18.31 99.49 $\pm$ 0.18 33.93 $\pm$ 2.51 32.91 $\pm$ 2.10 97.77 $\pm$ 4.22 96.77 $\pm$ 4.12 95.77 $\pm$ 4.02 98.77 $\pm$ 4.32 wine 82.08 $\pm$ 9.50 99.92 $\pm$ 0.12 100.00 $\pm$ 0.00 66.04 $\pm$ 11.09 70.33 $\pm$ 1.55 89.30 $\pm$ 2.14 73.75 $\pm$ 20.60 87.48 $\pm$ 0.21 71.24 $\pm$ 10.47 75.83 $\pm$ 1.79 77.89 $\pm$ 2.07 96.67 $\pm$ 3.78 90.00 $\pm$ 6.49 84.00 $\pm$ 1.11 83.00 $\pm$ 1.01 82.00 $\pm$ 0.91 85.00 $\pm$ 1.21 WPBC 49.53 $\pm$ 16.13 47.74 $\pm$ 2.20 51.33 $\pm$ 3.43 46.44 $\pm$ 3.00 56.34 $\pm$ 3.89 65.49 $\pm$ 1.65 47.17 $\pm$ 2.92 66.20 $\pm$ 1.21 59.52 $\pm$ 7.09 57.31 $\pm$ 3.88 70.38 $\pm$ 1.21 45.81 $\pm$ 6.76 58.06 $\pm$ 11.96 55.74 $\pm$ 5.49 54.74 $\pm$ 5.39 53.74 $\pm$ 5.29 56.74 $\pm$ 5.59 yeast 42.06 $\pm$ 2.20 42.67 $\pm$ 1.39 37.02 $\pm$ 3.11 49.48 $\pm$ 0.57 45.29 $\pm$ 0.42 54.13 $\pm$ 0.68 43.23 $\pm$ 3.54 68.79 $\pm$ 0.08 47.98 $\pm$ 10.23 42.41 $\pm$ 0.84 66.41 $\pm$ 1.18 38.50 $\pm$ 3.34 41.68 $\pm$ 3.77 53.01 $\pm$ 1.71 52.01 $\pm$ 1.61 51.01 $\pm$ 1.51 54.01 $\pm$ 1.81 breastw 98.72 $\pm$ 0.60 93.54 $\pm$ 1.39 99.09 $\pm$ 0.57 66.23 $\pm$ 2.81 51.18 $\pm$ 9.24 66.04 $\pm$ 2.92 99.06 $\pm$ 0.50 92.92 $\pm$ 0.83 83.10 $\pm$ 9.15 50.16 $\pm$ 7.61 90.40 $\pm$ 1.00 99.42 $\pm$ 0.27 99.17 $\pm$ 0.58 57.55 $\pm$ 13.03 56.55 $\pm$ 12.93 $5$$\pm$ 12.83 58.55 $\pm$ 13.13 campaign 71.14 $\pm$ 0.89 66.42 $\pm$ 0.21 72.19 $\pm$ 0.86 62.54 $\pm$ 0.83 69.26 $\pm$ 1.81 70.14 $\pm$ 0.99 77.42 $\pm$ 0.54 73.89 $\pm$ 0.40 59.40 $\pm$ 6.16 71.52 $\pm$ 2.12 69.38 $\pm$ 0.64 73.97 $\pm$ 0.33 75.07 $\pm$ 1.24 63.72 $\pm$ 3.61 62.72 $\pm$ 3.51 61.72 $\pm$ 3.41 64.72 $\pm$ 3.71 cardio 93.06 $\pm$ 1.79 93.04 $\pm$ 0.64 84.28 $\pm$ 2.89 81.24 $\pm$ 2.33 60.66 $\pm$ 8.68 97.71 $\pm$ 1.10 93.53 $\pm$ 0.69 99.40 $\pm$ 0.10 99.17 $\pm$ 0.13 38.83 $\pm$ 10.40 98.70 $\pm$ 0.16 95.41 $\pm$ 0.93 82.53 $\pm$ 2.40 96.80 $\pm$ 1.36 95.80 $\pm$ 1.26 94.80 $\pm$ 1.16 97.80 $\pm$ 1.46 Cardiotocography 72.66 $\pm$ 2.96 78.33 $\pm$ 1.17 62.76 $\pm$ 2.08 59.71 $\pm$ 1.64 84.87 $\pm$ 9.38 98.98 $\pm$ 0.42 80.09 $\pm$ 2.85 99.29 $\pm$ 0.14 98.04 $\pm$ 0.77 86.65 $\pm$ 8.37 98.94 $\pm$ 0.24 66.48 $\pm$ 4.13 49.87 $\pm$ 4.02 98.54 $\pm$ 1.13 97.54 $\pm$ 1.03 96.54 $\pm$ 0.93 99.54 $\pm$ 1.23 celeba 69.95 $\pm$ 1.93 68.50 $\pm$ 0.27 60.34 $\pm$ 0.86 70.20 $\pm$ 0.36 67.81 $\pm$ 1.65 39.59 $\pm$ 1.76 75.60 $\pm$ 0.68 34.49 $\pm$ 0.27 31.31 $\pm$ 9.90 66.59 $\pm$ 7.27 34.81 $\pm$ 2.83 80.54 $\pm$ 1.43 78.18 $\pm$ 0.55 67.90 $\pm$ 4.12 66.90 $\pm$ 4.02 65.90 $\pm$ 3.92 68.90 $\pm$ 4.22 census 61.88 $\pm$ 2.97 53.35 $\pm$ 0.15 64.49 $\pm$ 0.24 50.34 $\pm$ 0.10 44.66 $\pm$ 3.87 48.68 $\pm$ 3.14 66.00 $\pm$ 0.31 51.87 $\pm$ 3.09 50.07 $\pm$ 3.44 43.61 $\pm$ 3.13 54.81 $\pm$ 2.85 65.92 $\pm$ 0.39 65.74 $\pm$ 0.32 52.22 $\pm$ 3.47 51.22 $\pm$ 3.37 50.22 $\pm$ 3.27 53.22 $\pm$ 3.57 Average 76.35 71.86 73.56 66.36 60.13 74.76 74.12 72.56 63.90 59.78 73.81 75.97 77.36 73.21 72.31 71.41 74.12